OpenStudy (anonymous):

how do you find the domain and range for p(t)=2800+1500 when t=0

OpenStudy (anonymous):

p(t) does not appear to be a function of time.

OpenStudy (anonymous):

I didnt type the whole question....... its says: find the relevent domain and range for the following function: the function p(t)=2800+1500t computes the population of a city from the years 1908 to 2008 with t=0 representing 1980

OpenStudy (anonymous):

If t= 0 represents 1980, what does t= 28 represent?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Let me know...

OpenStudy (anonymous):

idk what you are aking. im so confused

OpenStudy (anonymous):

If t= 0 represents 1980, then t= 1 would represent 1981 etc so what does t= 28 represent?

OpenStudy (anonymous):

oh got ya! it is 2008..

OpenStudy (anonymous):

So u are not allowed to enter any t which is outside 1980 to 2008 ie from t= 0 to t= 28, that is your domain. Your range is all the numbers that are produced by entering all these t's. So when t is 0, p(t=0) is 2800+1500*0 = 2800 and when t is 28 p(t=28) =2800+1500*28 = 44,800. So your range is from 2800 to 44800.

OpenStudy (anonymous):

omg thank you so much. that helped a ton!

OpenStudy (anonymous):

ur welcome.

OpenStudy (anonymous):

the range was wrong.... instead of 44800 is is 70000....why?

OpenStudy (anonymous):

I don't know, let me look again...

OpenStudy (anonymous):

70000 is wrong.

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

The equation is 2800+1500t, right?

OpenStudy (anonymous):

hmm... well thats why my thing says... its for college algebra on my math lab

OpenStudy (anonymous):

And the maximum value for t is 28 so 2800 + 1500*28 = 44800.

OpenStudy (anonymous):

I notice your question says "the years 1908 to 2008 with t=0 representing 1980 Is 1908 a typo

OpenStudy (anonymous):

Must be or u would have negative population.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To be 70000 = 2800 + 1500t -> t= 44.8 which is obviously wrong.

OpenStudy (anonymous):

44800 is right.

OpenStudy (anonymous):

Srictly, from 2800 to 44800.

OpenStudy (anonymous):

Did it say the domain was right?

OpenStudy (anonymous):

yeah.... well thanks anyways

OpenStudy (anonymous):

You should complain....:-)

OpenStudy (anonymous):

i did same prob and diff num and got it right